If a number is increased by 21, the result is 30. What is the number?
Found 2 solutions by Abbey, CharStar:
Answer by Abbey(339)

let x = the number
x+21=30
subtract 21 from both sides:
x=9
the number is 9

Answer by CharStar(110)

Let x = a number - (1) number
x + 21 = 30
x + 21 = 30
Subtract 21 from both sides
x = 9
Check your answer
9 + 21 = 30